I was wondering if someone could tell me if I got these problems right? I am finding this to be a difficult subject.
A pin consists of 2 letters followed by 3 numbers (assume repetition is allowed and order is important).
a) how many different pin numbers are there?
26*26*10*10*10=676,000
b) what is the proabilty a pin number selected randomly ends in 000?
26*26*9*9*9=492,804
c)what is the probability a pin selected at random begins with AA?
25*25*10*10*10=625,000
d)what is the probability a pin selected at random being with A and ends with 00?
25*26*10*9*9=526,500
e)what is the probability a pin selected at random has a repeated letter and different numbers?
26*25*10*9*8=468,000

Thanks for your time and help,
Gina

Answer by stanbon(75887)

A pin consists of 2 letters followed by 3 numbers (assume repetition is allowed and order is important).
a) how many different pin numbers are there?
26*26*10*10*10=676,000
b) what is the probabilty a pin number selected randomly ends in 000?
26*26*9*9*9=492,804
So probability = 492,804/676,000
--------------------------------------
c)what is the probability a pin selected at random begins with AA?
25*25*10*10*10=625,000
So probability = 625,000/676,000
--------------------------------------
d)what is the probability a pin selected at random begins with A and ends with 00?
1*26*10*1*1=260
So probability = 260/676,000
--------------------------------------
e)what is the probability a pin selected at random has a repeated letter and different numbers?
That depends on the rules followed for building the pin.
If repetition is allowed in building the pin the probability is 1.
================
Cheers,
Stan H.
===========